OpenStudy (anonymous):

just need work cheek will have pic below and the work below too solve for x and y

OpenStudy (anonymous):

show us beautiful lol

OpenStudy (anonymous):

 x+105=80+2x x=25 20+4y+y=180 y=32

jimthompson5910 (jim_thompson5910):

both are correct

OpenStudy (anonymous):

kk thanks :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

agree with jim